# Radio-loudness statistics of quasars from Quaia–VLASS

### v1

@@ 2.4 Reddening correction
where $G^{\prime}$ is the uncorrected magnitude, $G$ is extinction-corrected magnitude, and $r$ is the reddening in the given region of the sky as provided by Schlegel et al. (1998). Further, $\langle R_{\zeta}\rangle\approx 2.74$ is the extinction coefficient for the Gaia $G$-band as provided by Casagrande & VandenBerg (2018). Except stated otherwise, all subsequent references in this paper to the $G$-band magnitude refer to the corrected value from Eq. 2.
@@ 3 Results
Using the cross-matched Quaia–VLASS sources, we estimated the radio loudness of our sample in various subsets and using different cuts. We used the definition by Ivezić et al. (2002) for the radio magnitude $t$ as
@@
where $F_{\mathrm{tot}}$ is the total flux density of the source (in the $S$-band). In Fig. 5, all the cross-matched Quaia–VLASS sources are displayed in a $G-t$ diagram, including a binning along the $G$ axis for future tests. Again following Ivezić et al. (2002), we used the $G-t$ magnitude difference as a measure for the radio loudness of the sources. We defined the following two cases:
$\begin{array}[]{ll}G-t>2.5:&\text{{\bf radio-loud sources}}\\ G-t\leq 2.5:&\text{radio-quiet sources}\end{array}$ (4)
Table 2: Differences between our work and Ivezić et al. (2002).
$\hline\cr\hline\cr Present work&\text{Previous work}\\ \hline\cr\emph{Quaia}--VLASS&\text{SDSS--FIRST}\\ $G$-band + $S$-band&\text{$i$-band + $L$-band}\\ 43,650 cross-matches&\text{$3066$ cross-matches}\\ \hline\cr\end{array}$
This definition was derived from the equivalent calculation in Ivezić et al. (2002), $R_{m}=0.4(m-t)$, where $m$ is a magnitude in the optical, $t$ is the radio magnitude at $1.4$ GHz (in that example, $R_{m}>1$ corresponds to radio-loud and $R_{m}\leq 1$ to radio-quiet objects).
@@
Figure 6: The radio-loud fraction $N_{\rm loud}/N_{\rm total}$ of QSOs in different bins as a function of $G$ (black curve) and the number $N_{\rm total}$ of Quaia quasars in these bins (blue curve). The colour-coding along $G$ is equivalent to the colour-coding in Fig. 5. The black curve shows that even though it is expected to underestimate the radio-loud fraction at $G>18.5$, such a trend of a sudden drop in the curve is not observed.
3.1 A high-purity catalogue of radio-loud quasars
In Fig. 5, a solid line depicts the border between radio-loud and radio-quiet sources, as expressed in Eq. 4. Sources are coloured according to bins in $G$, which correspond to optically bright, moderate, and faint bins (the number of sources within different $G$-bins is given in Table 1).
@@ 3.1 A high-purity catalogue of radio-loud quasars
The choice of $G$-band range values for these bins is motivated by a desire to keep these splits simple, yet offering meaningful comparisons to Ivezić et al. (2002). We note that while we kept $t$ and radio loudness consistent for the sake of comparison, there are some inevitable technical differences as displayed in Table 2. The main addition is a significantly larger sky coverage, a larger cross-matched sample, and thus a reduction of possible biases in the analysis.
@@
While the actual definitions of radio loudness may differ (as well as their range of validity), we note that of the 43,650 cross-matched Quaia–VLASS quasars shown in Fig. 5 the vast majority is clearly visible in the radio (41,496 of them are radio-loud, given our definition). Therefore, an important outcome of our work is a high-purity catalogue of quasars with radio counterparts, which may be used in various other projects in the future.
@@
Considering the completeness of our radio-loud catalogue (besides its purity), one should also consider the sensitivity limits of the input optical and radio data sets, that may bias the radio-loudness statistics for faint sources. We provide various tests of these systematic effects in the following sections, applying different binning schemes and cuts to the Quaia–VLASS sample to estimate possible variations in the radio-loud fraction.
@@
When estimating the radio-loud fraction ($N_{\rm loud}/N_{\rm total}$) of Quaia–VLASS sources, it is reasonable to restrict the sample to a range where the sensitivity of the instruments is reliable, i.e. where the measured $N_{\rm loud}$ and $N_{\rm total}$ source counts are complete within a chosen magnitude limit.
@@
Ivezić et al. (2002) applied such a restriction in the form of a cut in the optical $i$-band, providing a fiducial value for the fraction of radio-loud QSOs with $i<18.5$. This particular cut follows from their radio-loudness formula ($G-t>2.5$) and the sensitivity limit they work with in the radio ($t<16.0$, that is slightly more conservative than the nominal $t<16.5$ where they still find radio sources).
@@
Figure 7: The $G-t$ magnitude difference, i.e. our proxy for radio loudness, is averaged in various bins. Top: we bin our catalogue on the Quaia selection function (completeness) values, and also illustrate the role of the Galactic extinction correction. Bottom: we bin our sample on redshift, and also on $G$ magnitudes (see colours), and identify weak trends at the lowest redshifts $(z<0.5)$.
@@
In our work, we used the same Eq. 4 for defining radio loudness, and the VLASS has a similar sensitivity limit in $t$ (see Fig. 5). While this condition motivates a cut in the Quaia data at $G<18.5$, we tested the whole $G<20.5$ magnitude range to measure $N_{\rm loud}/N_{\rm total}$ for a comprehensive analysis.
@@
Figure 5 depicts the origin of this unreliability. For fainter Quaia magnitudes, the radio magnitudes may fall outside of the sensitivity limit $t\approx 16.0$, which suggests lower and lower completeness when one intends to estimate the radio-loud fraction on the faint end (while the purity of the sample is expected to remain high, as we argued above).
@@
In Fig. 6, we show how the radio-loud fraction of QSOs drops for fainter optical sources (which are naturally more abundant in the sample). We made the following observations:
  • The black line represents the radio-loud fraction of sources as a function of the $G$ magnitude, showing a linear drop in the radio-loudness fraction.
@@
  • Also, we observed that the radio-loud fraction is not constant even in the range of the brightest sources (e.g. $G<17$), where high completeness is expected in both $N_{\rm loud}$ and $N_{\rm total}$.
@@
  • Further, constraining the measurement of the radio-loud fraction to $G<18.5$ would ignore the bulk of the Quaia QSOs (see the blue curve in 6) and the majority of the cross-matched sources (see Table 1).
@@
For the above three reasons, we decided to express the radio-loud fraction of the Quaia–VLASS sources as a function of $G$ up to $G=20.5$, and let future users of our catalogue decide how they wish to balance purity and completeness.
@@
Figure 8: Radio-loud QSO fraction values in different parts of the sky are shown. We find significant fluctuations near the Milky Way, i.e. where the Quaia catalogue is knowingly less complete (although on average there are no clear biases near the Galactic plane either, see top panel of Fig. 7). In the cleanest areas of the sky, we do not identify any relevant anisotropy in the radio-loud fraction.
@@
Nevertheless, if we apply the measure for the radio-loud fraction by Ivezić et al. (2002), then we arrive at a value of $N_{\rm loud}/N_{\rm total}\approx 8.04\%\pm 0.09\%$ radio-loud fraction for $G<18.5$, while they estimate $N_{\rm loud}/N_{\rm total}\approx 8\%\pm 1\%$ for a similarly bright, conservatively selected subset of their data ($i<18.5$). Again, we note the differences between the actual bands used by Ivezić et al. (2002) and our analysis, but we conclude that consistent radio-loud fraction values are observed for the brightest QSOs.
3.3 Radio loudness vs. QSO completeness
@@
In Fig. 7, we show multiple tests that we performed by binning and modifying our data. In the context of the Quaia selection function (a measure for the completeness of QSO detections per pixel), we show on the top panel the role of the $G$-band extinction correction to $G-t$ (our proxy for the radio loudness), and the overall spread of the Quaia–VLASS catalogue as a function of this systematic parameter of the input data.
@@
The colour of the data points corresponds to their $G$-band magnitudes (as shown in Fig. 5). Empty circles mark the average value of $G^{\prime}-t$ within a given bin of the selection function values, considering the raw $G^{\prime}$ magnitudes (not corrected for extinction). In contrast, full black circles show the binned average $G-t$ values after extinction correction.
@@
From the comparison to the overall mean $G-t$ value of the whole Quaia–VLASS sample (horizontal line), we found that the extinction correction is effective in reducing the deviations between $G-t$ values measured at low vs. high selection function values. The top panel of Fig. 7 demonstrates that most filled circles fall within $1\sigma$ of the global mean $G-t$ value for all cross-matches. This result further justifies our use of the extinction-corrected $G$-band values in subsequent analyses, and it suggests that Galactic dust extinction will not significantly affect our radio-loudness analysis.
3.4 Radio loudness vs. redshift
@@
To perform a further test of systematics, we looked for possible changes in radio loudness as a function of QSO redshift. As shown in Fig. 7 (bottom panel), we observed a moderately significant bias in the average $G-t$ distribution (i.e. in the proxy for radio loudness) at the lowest redshifts ($z<0.5$). We estimated this statistic for the three $G$-bins that we considered above, and we found almost identical trends as for the $G<20.5$ case.
@@
The source of this biased $G-t$ distribution at the low-redshift part of the Quaia QSO catalogue remains unclear. One possibility is that the star–galaxy–QSO separation in the GaiaWISE colour space is less efficient, or the photometric redshifts of the quasars might also be biased at the lowest redshifts where the training set based on Sloan Digital Sky Survey (SDSS) data is less complete. For a detailed explanation, further measurements and statistical probes would be required, which are beyond the scope of this paper.
@@
Figure 9: Radio-loud fraction of Quaia sources colour-coded as a function of the absolute magnitude $G_{\rm abs}$ and the redshift $z$. We observed significant noise due to small-number statistics at the bright-end pixels, leading to strong fluctuations in the radio-loudness fraction.
3.5 Radio-loud fraction across the sky
@@
Looking for possible large-scale spatial patterns on the sky, we created a pixelized map of the radio-loud QSO fraction within each pixel. In Fig. 8, a rather homogeneous distribution is shown across the full observed area, except close to the Galactic plane.
@@
We note that the fluctuations in radio loudness are naturally stronger at low Galactic latitudes, due to sparsity of the QSOs (see the top panel of Fig. 1). For a conservative analysis, future users of our Quaia–VLASS catalogue may apply specific cuts in the Quaia selection function map [^2] Publicly available at https://zenodo.org/records/10403370, offering convenient ways of limiting the analyses to the cleanest and most complete parts of the sky.
@@
Additionally, we tested the possible dependence of the radio-loud fraction on the absolute magnitudes of the quasars. Following Ivezić et al. (2002), we used the absolute magnitude $G_{\rm abs}$, which we calculated as
@@
where $d_{\rm lm}(z)$ is the luminosity distance [^3] Assuming a Planck $\Lambda$CDM cosmology (Planck Collaboration et al., 2020) of the object, calculated with astropy [^4] https://www.astropy.org/.
@@
Given the above caveats, our main findings regarding the radio-loud fraction patterns in $G_{\rm abs}$ and redshift bins are illustrated in Fig. 9. Each pixel displays the radio-loud fraction of Quaia sources in that bin, using a colour-coding scheme to easily identify any possible trend in the data along either axis.
The figure allows a separate analysis of constant $z$ values in rows against varying $G_{\rm abs}$ values in columns, or vice versa. We found that for a fixed redshift value, radio-loud fraction increases towards brighter $G_{\rm abs}$ bins. Similarly, for a fixed $G_{\rm abs}$, the binned radio-loud fraction values show a decreasing trend with increasing redshift.
We note that, besides the known redshift errors, our $G_{\rm abs}-z$ plot in Fig. 9 displays significant noise due to the sparsity of objects in pixels with the most negative $G_{\rm abs}$ values, leading to large fluctuations in radio-loud fractions. Hence we suggest that QSOs falling in these pixels should be flagged or ignored in any subsequent analyses of redshift- and magnitude-dependence of radio-loud fraction that use our Quaia–VLASS catalogue.
Figure 10: Normalized histograms of the radio-loud and radio-quiet populations as a function of $G_{\rm abs}$. On the left, we show the full sample of cross-matches, while on the right a $G<18.5$ cut is applied.

### v2

@@ 2.4 Reddening Correction
where $G^{\prime}$ is the uncorrected magnitude, $G$ is the extinction-corrected magnitude, and $r$ is the reddening in the given region of the sky as provided by Schlegel et al. (1998). Further, $\langle R_{\zeta}\rangle\approx 2.74$ is the extinction coefficient for the Gaia $G$-band as provided by Casagrande & VandenBerg (2018). Except stated otherwise, all subsequent references in this paper to the $G$-band magnitude refer to the corrected value from Eq. 2.
@@ 3 Results
Using the cross-matched Quaia–VLASS sources, we estimated the radio loudness of our sample via the definition by Ivezić et al. (2002) for the radio magnitude $t$ as
@@
where $F_{\mathrm{tot}}$ is the total flux density of the source (in the $S$-band). Within the same work the radio loudness $R_{m}$ is expressed as the difference between the optical and radio magnitudes of the sources: $R_{m}=0.4(m-t)$, where $m$ is a magnitude in the optical, $t$ is the radio magnitude at $1.4$ GHz (in that example, $R_{m}>1$ corresponds to radio-loud and $R_{m}\leq 1$ to radio-quiet objects). The equation for $R_{m}$ turned into an inequality translates in our work to the following definitions, used throughout this paper:
[…]
3.1 G-t color space and histogram
The distribution of radio-loud and radio-quiet cross-matches in a GtG-t color space is shown in Figure 4. Sources are colored according to bins in $G$, which correspond to optically bright, moderate, and faint bins (the number of sources within different $G$-bins is given in Table 1). The choice of $G$-band range values for these bins is motivated by a desire to keep these splits simple, yet offering meaningful comparisons to Ivezić et al. (2002). Visually, the sensitivity limit in the radio lies within $16<t_{\mathrm{lim}}<17$. This limit can also be calculated via the conventional assumption of being equal 5 times the root mean square (RMS) noise limit (Arnaudova et al., 2024). Lacy et al. (2022) provide the flux density restricted by RMS noise as $\sigma_{\mathrm{fd}}=170\,\mu\mathrm{Jy}$. Inserting $F_{\mathrm{tot}}=5\sigma_{\mathrm{fd}}$ in equation 3 gives a magnitude limit of $t\approx 16.6$, which we plotted in Figure 4 as a horizontal dashed line and it agrees well with the visual extent of the data points.
We treat all Quaia sources, for which we do not find a matching radio counterpart according to the methodology in Section 2.3, as radio-quiet QSOs. All RLF values provided throughout this work are based on this assumption. In that way we make use of all the available optical sources, even if they do not have measured radio magnitudes. This approach provides accurate RLF values for the sources with $G\lessapprox 18.5$, because within this regime the radio loudness border at Gt=2.5G-t=2.5 is above the VLASS sensitivity limit (see Figure 4, for all blue data points the solid line is above the dashed line). On the other hand, for sources with $G\gtrapprox 18.5$, our approach of treating radio-undetected sources as radio-quiet will lead to an underestimation of the RLF - in this regime Quaia sources could have radio-counterparts fulfilling Gt>2.5G-t>2.5 that happen to have $t$ fainter than the sensitivity limit of VLASS (see the indeterminate region in Figure 4).
An alternative approach to treating radio-undetected sources is presented in Arnaudova et al. (2024). There, the radio-undetected sources get assigned a radio magnitude equal to 5 times the local RMS value taken from the radio-survey RMS maps at the coordinates of the quasars. In the present work, this approach would result in a mean sensitivity limit in the radio of $t\approx 16.6$. This method would lead to an overestimation of the RLF at faint $G$ – expressing this in Figure 4 would mean to populate all radio-undetected sources along the dashed line, hence the optically faintest will occupy the indeterminate region and be radio-loud. Nonetheless, this approach has the advantage of enabling a GtG-t histogram for all sources, including the radio-undetected ones. We show this histogram of our data in Figure 5. It allows for naive check of the quasar dichotomy hypothesis – one observes that both the radio-undetected Quaia sources (blue) and the radio-detected ones (red and orange) have smooth distributions without strong peaks indicative for subclasses in the quasar population. This is consistent with results from other recent works, reporting the transition from radio-quiet to radio-loud QSO being smooth and lacking evidence of a clear dichotomy (Gürkan et al., 2019; Macfarlane et al., 2021; Arnaudova et al., 2024). Nonetheless, interpretations of Figure 5 are to be done with caution, since data are limited and $t$ values for the blue histogram are only upper limits. In Section 4, we point out future surveys that may provide new insight on these patterns by using even better data sets.
Figure 5: The distribution of the radio loudness measure GtG-t for all Quaia sources and the Quaia–VLASS cross-matches. The vertical black line is the delineation between radio-quiet ($G-t\leq 2.5$) and radio-loud (Gt>2.5G-t>2.5) sources. The blue arrow indicates that the shape of the blue distribution would shift to the left, if for these sources real measurements for $t$, instead of estimates, were present.
@@ 3.1 G-t color space and histogram
Figure 6: The RLF $N_{\rm loud}/N_{\rm quaia}$ of QSOs in different bins as a function of $G$ (black curve), the number of Quaia quasars within these bins (blue histogram) and the number of radio-loud crossmatches (orange histogram). The background color-coding along $G$ is equivalent to the color-coding in Figure 4. The black curve shows that even though it is expected to underestimate the RLF at $G>18.5$ due to the effect of the indeterminate region (see Figure 4), such a trend of a sudden drop in the curve is not observed. The errors are calculated via the binomial standard error $\sqrt{f(1-f)/n_{\rm b}}$, where $f$ is the RLF in a bin and $n_{\rm b}$ the total number of QSOs in the bin.
3.2 A High-purity Catalog of Radio-loud Quasars
@@
In the above Section, we kept the definitions for $t$ (Equation 3) and radio loudness (Equation 4) consistent with Ivezić et al. (2002) for the sake of comparison. Nonetheless, there are some inevitable technical differences, as displayed in Table 2. The main contribution of this work is a significantly larger sky coverage, a larger cross-matched sample, and thus a reduction of possible biases in the analysis.
@@
While the actual definitions of radio loudness may differ (as well as their range of validity), we note that of the 43,650 cross-matched Quaia–VLASS quasars shown in Figure 4 the vast majority is clearly visible in the radio (41,496 of them are radio-loud, given our definition). Therefore, an important outcome of our work is a high-purity catalog of quasars with radio counterparts, which may be used in various other projects in the future.
@@
Considering the completeness of our radio-loud catalog (besides its purity), one should also consider the sensitivity limits of the input optical and radio data sets, that may bias the radio loudness statistics for faint sources. We provide various tests of these systematic effects in the following sections, applying different binning schemes and cuts to the Quaia–VLASS sample to estimate possible variations in the RLF.
3.3 Comparison to Previous Works
As evident from Table 2, row “$N$ cross-matches”, by using the same methodology as Ivezić et al. (2002) we discover $\approx 14$ times more optical–radio QSO cross-matches. The same table shows that our work is superior to the other compared works in terms of cross-matches due to the larger radio and optical parent catalogs we used. Many of the previous works have used SDSS QSO samples containing fainter sources than in our work. Nonetheless, all sources we observe outside the SDSS footprint are novel to the compared previous works. This means, by simply comparing the survey overlap areas in Table 2 and making the reasonable approximation that the Quaia–VLASS sample is homogenous, our subset of optical-radio cross-matches outside the sky coverage of Baloković et al. (2012) is $\approx 32,000$ ($\approx 30,000$ of them radio-loud). Likewise, our subset of cross-matches outside of the Arnaudova et al. (2024) sky area is $\approx 35,000$ ($\approx 33,000$ of them radio-loud). By covering a significantly larger sky area, our catalog is superior to the compared samples, for users desiring the closest possible representation of the full sky at present. Also, in contrast to some of the previous works, our code and final data products are freely available to the public, see Section 5. Our catalog is also superior for cosmological applications, since the Gaia scanning pattern is more homogenous than selective SDSS QSO observations and the Quaia selection function can help further mitigate selection biases. On the other hand, the SDSS spectroscopic observations are superior to the photometric ones by Gaia, so redshift and classification accuracy are better in the previous studies in Table 2.
Table 2: Comparison of Quasar radio loudness Studies
Parameter Ivezić et al. (2002) Kimball & Ivezić (2008) Baloković et al. (2012) Arnaudova et al. (2024) this work
Cross-match area [deg2] 1,230 2,894 7,600 5,634 29,230
NN optical sourcesa 25.3M (all object types) 35,450 98,544 189,680 1.085M
NN radio sourcesb 107,654 140,000 4.4M (in full LoTSS area) 1.874M
NN cross-matches 3,066 2,927 8307 33,968 43,650
NN cross-matches, loud 3,697c 41,496
Optical survey (band) SDSS ($i$, $r$) SDSS DR5 ($i$, $g$, $r$) SDSS DR7 ($i$) SDSS DR14 ($i$) Gaia ($G$)
Optical limit $i<21$ $i\lesssim 21$ $i\lesssim 21.2$ $i\lesssim 21$ $G<20.5$
Radio survey (frequency) FIRST (1.4 GHz) NVSS, FIRST (both 1.4 GHz) FIRST (1.4 GHz) LoTSS DR2 (144 MHz) VLASS2 QL (3 GHz)
Radio limitd 1 mJy ($t\approx 16.4$) 1 mJy ($t\approx 16.4$) 1 mJy ($t\approx 16.4$) 0.415 mJy ($t\approx 17.35$) 0.85 mJy ($t\approx 16.6$)
@@ 3.3 Comparison to Previous Works
When estimating the RLF ($N_{\rm loud}/N_{\rm total}$) of Quaia–VLASS sources, it is reasonable to restrict the sample to a range where the sensitivity of the instruments is reliable, i.e. where the measured $N_{\rm loud}$ and $N_{\rm total}$ source counts are complete within a chosen magnitude limit.
@@
Ivezić et al. (2002) applied such a restriction in the form of a cut in the optical $i$-band, providing a fiducial value for the fraction of radio-loud QSOs with $i<18.5$. This particular cut follows from their radio loudness formula (Gt>2.5G-t>2.5) and the sensitivity limit they work with in the radio ($t<16.0$, that is slightly more conservative than the nominal $t<16.5$ where they still find radio sources).
@@
Figure 7: The GtG-t magnitude difference, i.e. our proxy for radio loudness, and its median in various bins. All error bars are equal to $\mathrm{median}(G-t)/\sqrt{n_{\rm c}}$ with $n_{\rm c}$ being the number of cross-matched QSOs in this bin and $\mathrm{median}(G-t)$ being the median within the same bin. Top: we bin our catalog on the Quaia selection function (completeness) values, and also illustrate the role of the Galactic extinction correction. Bottom: we bin our sample on redshift, and also on $G$ magnitudes (see colors), and identify weak trends at the lowest redshifts $(z<0.5)$.
@@
In our work, we used the same Eq. 4 for defining radio loudness, and the VLASS has a similar sensitivity limit in $t$ (see Figure 4). While this condition motivates a cut in the Quaia data at $G<18.5$, we tested the whole $G<20.5$ magnitude range to measure $N_{\rm loud}/N_{\rm total}$ for a comprehensive analysis.
@@
Figure 4 depicts the origin of this unreliability. For fainter Quaia magnitudes, the radio magnitudes may fall outside of the sensitivity limit $t\approx 16.0$, which suggests lower and lower completeness when one intends to estimate the RLF on the faint end (while the purity of the sample is expected to remain high, as we argued above).
In Figure 6, we show the RLF of QSOs as a function of $G$ (denoted by the black line), where we make the following observations:
@@
  • The RLF of sources drops linearly for fainter optical sources (which are naturally more abundant in the sample).
@@
  • the RLF is not constant even in the range of the brightest sources (e.g. $G<17$), where high completeness is expected in both $N_{\rm loud}$ and $N_{\rm total}$.
@@
  • Constraining the measurement of the RLF to $G<18.5$ would ignore the bulk of the Quaia QSOs (see the blue histogram in 6) and the majority of the cross-matched sources (see Table 1).
@@
For the above three reasons, we decided to express the RLF of the Quaia–VLASS sources as a function of $G$ up to $G=20.5$, and let future users of our catalog decide how they wish to balance purity and completeness.
@@
Figure 8: Quasar RLF in different parts of the sky. We find significant fluctuations near the Milky Way, i.e. where the Quaia catalog is knowingly less complete (although on average there are no clear biases near the Galactic plane either, see top panel of Figure 7). In the cleanest areas of the sky, we do not identify any relevant anisotropy in the RLF.
@@
Nevertheless, if we apply the measure for the RLF by Ivezić et al. (2002), then we arrive at a value of $N_{\rm loud}/N_{\rm total}\approx 8.04\%\pm 0.09\%$ RLF for $G<18.5$, while they estimate $N_{\rm loud}/N_{\rm total}\approx 8\%\pm 1\%$ for a similarly bright, conservatively selected subset of their data ($i<18.5$). Again, we note the differences between the actual bands used by Ivezić et al. (2002) and our analysis, but we conclude that consistent RLF values are observed for the brightest QSOs.
3.5 Radio Loudness vs. QSO Completeness, Redshift, and Sky Coverage
@@
In Figure 7, we show multiple tests that we performed by binning and modifying our data. In the context of the Quaia selection function (a measure for the completeness of QSO detections per pixel), we show on the top panel the role of the $G$-band extinction correction to GtG-t (our proxy for the radio loudness), and the overall spread of the Quaia–VLASS catalog as a function of this systematic parameter of the input data.
@@
The color of the data points corresponds to their $G$-band magnitudes (as shown in Figure 4). Empty circles mark the median value of $G^{\prime}-t$ within a given bin of the selection function values, considering the raw $G^{\prime}$ magnitudes (not corrected for extinction). In contrast, full black circles show the binned median GtG-t values after extinction correction.
@@
From the comparison to the median GtG-t value of the whole Quaia–VLASS sample (horizontal line in top of Figure 7), we found that the extinction correction is effective in reducing the deviations between GtG-t values measured at low vs. high selection function values. The top panel of Figure 7 demonstrates that most filled circles fall within $1\sigma$ of the global median GtG-t value for all cross-matches. This result further justifies our use of the extinction-corrected $G$-band values in subsequent analyses, and it suggests that Galactic dust extinction will not significantly affect our radio loudness analysis.
@@
To perform a further test of systematics, we looked for possible changes in radio loudness as a function of QSO redshift. As shown in Figure 7 (bottom panel), we observed a moderately significant bias in the median GtG-t distribution (i.e. in the proxy for radio loudness) at the lowest redshifts ($z<0.5$). We estimated this statistic for the three bins in G magnitude that we considered above, and we found almost identical trends as for the $G<20.5$ case.
@@
The source of this biased GtG-t distribution at the low-redshift part of the Quaia QSO catalog remains unclear. One possibility is that the star–galaxy–QSO separation in the Gaia–WISE color space is less efficient, or the photometric redshifts of the quasars might also be biased at the lowest redshifts where the training set based on Sloan Digital Sky Survey (SDSS) data is less complete. For a detailed explanation, further measurements and statistical probes would be required, which are beyond the scope of this paper.
@@
Figure 9: RLF of Quaia sources color-coded as a function of the absolute magnitude $G_{\rm abs}$ and the redshift $z$. We observed significant noise due to small-number statistics at the bright-end pixels, leading to strong fluctuations in the RLF. The inset shows the spread of $\Delta G_{\mathrm{abs}}$ being comparatively small to the grid-width of $G_{\mathrm{abs}}$ in the large figure.
@@
Looking for possible large-scale spatial patterns on the sky, we created a pixelized map of the radio-loud QSO fraction within each pixel. In Figure 8, a rather homogeneous distribution is shown across the full observed area, except close to the Galactic plane.
@@
We note that the fluctuations in radio loudness are naturally stronger at low Galactic latitudes, due to sparsity of the QSOs (see left side of Figure 1). For a conservative analysis, future users of our Quaia–VLASS catalog may apply specific cuts in the Quaia selection function map, offering convenient ways of limiting the analyses to the cleanest and most complete parts of the sky.
@@
Additionally, we tested the possible dependence of the RLF on the absolute magnitudes of the quasars. Following Ivezić et al. (2002), we used the absolute magnitude $G_{\rm abs}$, which we calculated as
@@
where $d_{\rm lm}(z)$ is the luminosity distance of the object.
@@
Given the above caveats, our main findings regarding the RLF patterns in $G_{\rm abs}$ and redshift bins are illustrated in Figure 9. Each pixel displays the RLF of Quaia sources in that bin, using a color-coding scheme to easily identify any possible trend in the data along either axis.
The figure allows a separate analysis of constant $z$ values in rows against varying $G_{\rm abs}$ values in columns, or vice versa. We found that for a fixed redshift value, the RLF increases towards brighter $G_{\rm abs}$ bins. Similarly, for a fixed $G_{\rm abs}$, the binned RLF values show a decreasing trend with increasing redshift. Applying a partial Spearman correlation to calculate the significance of the observed trends yields a correlation of the RLF with redshift (controlling for $G_{\rm abs}$) with values $r\approx-0.80$ and $p\approx 2.07\times 10^{-52}$ and a correlation of the RLF with $G_{\rm abs}$ (controlling for redshift) with values $r\approx-0.83$ and $p\approx 1.48\times 10^{-58}$. Hence, we can rule out that the trends are caused due to chance noise.
Below we discuss how much redshift errors influence the trends in Figure 9. The redshift error stated by Storey-Fisher et al. (2024) for the full Quaia catalog is $\mathrm{mean(|\Delta z/(1+z)|)\approx 0.01}$, significantly smaller than the grid size along $z$ in Figure 9. Storey-Fisher et al. (2024) assign the final redshift $z$ of a source, based on conditions, between $z=z_{kNN}$ (the redshift they estimate from a kNN method) and $z=z_{\rm Gaia}$ (the raw redshift for Quaia coming from Delchambre et al. (2023)). We follow the same conditions to pick for each source its corresponding redshift error standard deviation $\sigma_{z}$ from the Quaia catalog: $\sigma_{z}=$redshift_spz_err, if $z=z_{kNN}$, and $\sigma_{z}=$redshift_qsoc_upper, if $z=z_{\rm Gaia}$. Using these standard deviations, we assign to each source in Quaia probabilistically a new, erroneous redshift $z_{*}$ and corresponding $G_{\rm abs}(z_{*})$, and then calculate $\Delta G_{\rm abs}=G_{\rm abs}-G_{\rm abs}(z_{*})$, the fluctuation in absolute magnitude due to redshift uncertainties. $\Delta G_{\rm abs}$ is plotted as inset in Figure 9 and one can see that its spread is small compared to the $G_{\rm abs}$ grid size of 0.5 magnitudes in the same figure, and so are $\mathrm{median}(|\Delta G_{\rm abs}|)\approx 0.02$ and $\mathrm{mean}({|\Delta G_{\rm abs}|})\approx 0.17$. Only $\approx 17.3\%$ of the sources have a $\Delta G_{\rm abs}$ being outside of the 0.5 magnitude grid size width drawn with vertical lines in the inset plot. Hence, we conclude that redshift errors and their resulting $G_{\rm abs}$ errors are too small, compared to the grid size in both dimensions of Figure 9, to significantly influence trends observable in the figure.
Previous works report controversial results for the relations of RLF with luminosity and redshift. Some authors find no (or weak) correlation between RLF and both luminosity and redshift (Stern et al., 2000; Ivezić et al., 2002; Liu et al., 2021; Bañados et al., 2015; Keller et al., 2024). The latter two studies discuss redshift only. According to other researchers, RLF increases with decreasing redshift and increasing optical/UV luminosity (Jiang et al., 2007; Kratzer & Richards, 2015; Rusinek-Abarca & Sikora, 2021; Lah et al., 2023). In general, RLF can be influenced by the choice of: radio and optical bands; radio and optical flux limits; the exact value of the radio loudness threshold parameter; AGN variability, which affects mostly objects with radio loudness close to the demarcation value. The observed dependence of RLF on redshift could be explained e.g. in terms of the Mixture Evolution Scenario of the AGN Radio Luminosity Function (Yuan et al., 2017).
We note that, besides the known redshift errors, our $G_{\rm abs}-z$ plot in Figure 9 displays significant noise due to the sparsity of objects in pixels with the most negative $G_{\rm abs}$ values, leading to large fluctuations in RLFs. Hence, we suggest that QSOs falling in these pixels should be flagged or ignored in any subsequent analyses of redshift- and magnitude-dependence of RLF that use our Quaia–VLASS catalog.
Figure 10: Normalized histograms of the radio-loud and radio-quiet populations as a function of $G_{\rm abs}$. On the left, we show the full sample of cross-matches, while on the right a $G<18.5$ cut is applied. The error bars in the figure are Poisson errors equal to $n_{\rm norm}/\sqrt{n_{0}}$ with $n_{\rm norm}$ being the displayed normalized counts within the bin and $n_{0}$ the non-normalized counts in the same bin.